# Evaluating the Fitness of Ontologies for the Task of Question Generation

### v1

@@ 5 Conclusion
In this paper, the goal was to identify metrics for evaluating the fitness of ontologies for the task of AQG. The ROMEO methodology, a structured approach for identifying metrics based on specific task requirements was employed as a methodological framework. An expert-based metric derivation process was used to assess how different ontologies perform in AQG tasks. It was found that the characteristics of ontologies significantly influence the effectiveness of the question generation process. Our findings underscore that different ontologies yield varying levels of performance, highlighting the critical need to assess ontology quality in AQG. A set of structural and knowledge-based metrics was identified for assessing the fitness of ontologies for AQG tasks. This foundational study opens avenues for future work which should focus on exploring additional metrics that consider the content and user-oriented aspects of ontologies.

### v2

@@ 6 Conclusion
In this paper, we address the need for a way to characterise ontologies for use in ontology-based Automatic Question Generation. We propose a set of metrics, identified by using the ROMEO methodology that defines metrics based on specific task requirements. By consulting an expert in pedagogy, we defined a set of requirements and associated metrics to assess how different ontologies perform in AQG tasks. The metrics were then validated across an independent set of ontologies and their associated knowledge bases used in previous ontology-based AQG studies. Our results demonstrate that the characteristics of ontologies significantly influence the effectiveness of the question generation process, and that these characteristics are reflected accurately by these metrics. Our findings underscore that different ontologies yield varying levels of performance, highlighting the critical need to assess ontology quality in AQG.